OpenStudy (anonymous):

Henry grabs one shoe at a time at random without replacement from a box of ten shoes in his closet. The box contains 2 brown shoes, 4 black shoes, and 4 sneakers. What is the probability that the first two shoes he chooses will be brown? Round to the nearest percent if necessary. two percent twenty percent twenty-two percent forty percent

OpenStudy (anonymous):

what color are the sneakers?

OpenStudy (anonymous):

i dont know it doesnt say.

OpenStudy (anonymous):

then lets assume they are not brown

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

probability that the first two are brown is therefore \[\frac{2}{10}\times \frac{1}{9}\]

OpenStudy (anonymous):

im confused so whats the answer

OpenStudy (anonymous):

the answer i wrote above should be correct

OpenStudy (anonymous):

if you want a decimal, compute with a calculator

OpenStudy (anonymous):

so the answer would be 22%